# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v2

@@ 4.2.1 Bidirectional State Trajectory Generation
We utilize the pretrained forward and backward trajectory diffusion model in section 4.1 to generate bidirectional state trajectories. Prior methods [30, 27] model only the conditional distribution of trajectories originating from a given state, neglecting the historical pathways leading to those states. As a result, these data augmentation techniques yield only limited expansion in behavior diversity. To address this limitation without introducing additional noise, our proposed method, BiTrajDiff, generates both forward and backward trajectories and integrates them by stitching at a shared intermediate state $s_{t}$, which serves as a consistent anchor.
@@
$\hat{\tau}_{s}(s_{t})$ $=\left\{\mathbf{x}^{b}_{0}(\tau)[0:H-1],s_{t},\mathbf{x}^{f}_{0}(\tau)[1:H]\right\}$ (5)
$=\left\{\hat{s}_{t-H+1},\cdots,\hat{s}_{t-1},s_{t},\hat{s}_{t+1},\cdots,\hat{s}_{t+H-1}\right\},$ (6)
where $H$ is the horizon, and $\hat{s}_{i}$ is the diffusion generated states at MDP timestep ii. By conditioning on and concatenating at the intermediate anchor state $s_{t}$, we construct a state trajectory, $\hat{\tau}_{s}(s_{t})$, that connects $\hat{s}_{t-H+1}$ and $\hat{s}_{t+H-1}$. This process enables the connection of states that may be disconnected in the original dataset, thereby substantially enriching the diversity of observed behavioral patterns.
@@ 4.2.2 Bidirectional Trajectory Completion
Thus, by feeding the generated state trajectory $\hat{\tau}_{s}(s_{t})$ to the trained IDM and RM sequentially, we obtain the completed bidirectional trajectory: $\hat{\tau}(s_{t})\!=\!\left\{(\hat{s}_{i},\hat{a}_{i},\hat{r}_{i},\hat{s}_{i+1})\right\}_{i=t-H+1}^{t+H-2}$ [^1] For the sake of clearity, the condition state $s_{t}$ is writen as $\hat{s}_{t}$., where $\hat{a}_{i}=f_{\psi}(\hat{s}_{i},\hat{s}_{i+1})$ and $\hat{r}_{i}=r_{\phi}(\hat{s}_{i},\hat{a}_{i})$. Therefore, we can obtain a generated dataset $\mathcal{\hat{D}}=\{\hat{\tau}(s_{t_{i}})\}_{i=1}^{\hat{N}}$, where $\hat{N}$ is the number of generate trajectories.
@@ 4.2.3 Bidirectional Trajectory Filter
Greedy Trajectory Filter
The greedy trajectory filter selects the generated trajectories with the highest cumulative reward. Specifically, we sort the $C_{\text{ood}}$ trajectories retained from OOD trajectory filter by the sum reward $R(\bar{\tau})=\sum_{i=t-H+1}^{t+H-2}\hat{r}_{i}$. The top-$C_{\text{greedy}}$ trajectories with the highest sum rewards are picked as the final generated trajectories dataset $\tilde{\mathcal{D}}$ for offline RL training.
In summary, our BiTrajDiff operates in two key stages. Initially, it trains a bidirectional diffusion model to capture the distributions of both forward-future and backward-history trajectories. Subsequently, this trained model generates bidirectional trajectory sequences, all conditioned on intermediate anchor states. These generated sequences undergo further completion and filtering to create a high-fidelity augmented dataset. This augmented data is then integrated with the original datasets to train downstream offline RL algorithms. In our implementation, we iteratively sample batches of intermediate anchor states from the offline dataset $\mathcal{D}$, and generate bidirectional trajectories conditioned on these states using the aforementioned pipeline. The resulting trajectories are then incorporated into the final augmented dataset $\tilde{\mathcal{D}}$. We adopt CleanDiffuser [7] as the base backbone for BiTrajDiff, owing to its generality and robustness.
@@ 5 Experiments
To demonstrate the effectiveness of the proposed BiTrajDiff method, we conduct experiments on the D4RL benchmark [9]. Our evaluation seeks to answer the following questions: (1) Does BiTrajDiff outperform existing DA methods across various offline RL baselines? (Section 5.2) (2) Can BiTrajDiff generate more diverse trajectories compared to single-direction DA approaches? (Section 5.3)
@@ 5.1 Experiment Settings
Tasks and Datasets We conduct experiments on the D4RL benchmark [9], focusing on the Gym locomotion tasks. Specifically, we evaluate the performance of BiTrajDiff on three environments in MuJoCo tasks: HalfCheetah, Hopper, and Walker2d. Each environment includes three dataset types: medium, medium-replay, and medium-expert. The medium and medium-replay are collected from interactions of an unconverged policy with the environment, while the medium-expert dataset comprises a mixture of expert demonstrations and suboptimal trajectories.
Baselines and Evaluation Algorithms We compare our method with three state-of-the-art diffusion-based DA methods: Synther [30],DiffStitch [26] and RTDiff [47]. Meanwhile, to evaluate the effectiveness of the augmented trajectories of each DA method, we conduct extensive experiments utilizing two representative offline RL algorithms as base learners: IQL [23] and TD3+BC [10]. Both of the methods have been widely adopted as standard baselines in recent offline RL literature [36] due to their stable performance.
Table 1: Test returns of our proposed BiTrajDiff and baselines on the Gym tasks. $\pm$ corresponds to the standard deviation of the performance on 5 random seeds. The best and the second-best results of each setting are marked as bold and underline, respectively.
Source Task  IQL [23]  TD3BC [10]

### v3

@@ 4.2.1 Bidirectional State Trajectory Generation
We utilize the pretrained forward and backward trajectory diffusion model in section 4.1 to generate bidirectional state trajectories. Prior methods (lu2023synthetic, li2024augmenting) model only the conditional distribution of trajectories originating from a given state, neglecting the historical pathways leading to those states. As a result, these data augmentation techniques yield only limited expansion in behavior diversity. To address this limitation without introducing additional noise, our proposed method, BiTrajDiff, generates both forward and backward trajectories and integrates them by stitching at a shared intermediate state $s_{t}$, which serves as a consistent anchor.
@@
$\hat{\tau}_{s}(s_{t})=\left\{\mathbf{x}^{b}_{0}(\tau)[0\!:\!H\!-\!1],s_{t},\mathbf{x}^{f}_{0}(\tau)[1\!:\!H]\right\}=\left\{\hat{s}_{t-H+1},\cdots,\hat{s}_{t-1},s_{t},\hat{s}_{t+1},\cdots,\hat{s}_{t+H-1}\right\},$ (5)
where $H$ is the horizon, and $\hat{s}_{i}$ is the diffusion generated states at MDP timestep $i$. By conditioning on and concatenating at the intermediate anchor state $s_{t}$, we construct a state trajectory, $\hat{\tau}_{s}(s_{t})$, that connects $\hat{s}_{t-H+1}$ and $\hat{s}_{t+H-1}$. This process enables the connection of states that may be disconnected in the original dataset, thereby substantially enriching the diversity of observed behavioral patterns.
@@ 4.2.2 Bidirectional Trajectory Completion
Thus, by feeding the generated state trajectory $\hat{\tau}_{s}(s_{t})$ to the trained IDM and RM sequentially, we obtain the completed bidirectional trajectory: $\hat{\tau}(s_{t})\!=\!\left\{(\hat{s}_{i},\hat{a}_{i},\hat{r}_{i},\hat{s}_{i+1})\right\}_{i=t-H+1}^{t+H-2}***$For the sake of clarity, the condition state $s_{t}$ is written as $\hat{s}_{t}$., where $\hat{a}_{i}=f_{\psi}(\hat{s}_{i},\hat{s}_{i+1})$ and $\hat{r}_{i}=r_{\phi}(\hat{s}_{i},\hat{a}_{i})$. Therefore, we can obtain a generated dataset $\mathcal{\hat{D}}=\{\hat{\tau}(s_{t_{i}})\}_{i=1}^{\hat{N}}$, where $\hat{N}$ is the number of generate trajectories.
@@ 4.2.3 Bidirectional Trajectory Filter
Greedy Trajectory Filter. The greedy trajectory filter selects the generated trajectories with the highest cumulative reward. Specifically, we sort the $C_{\text{ood}}$ trajectories retained from OOD trajectory filter by the sum reward $R(\hat{\tau})=\sum_{i=t-H+1}^{t+H-2}\hat{r}_{i}$. The top-$C_{\text{greedy}}$ trajectories with the highest sum rewards are picked as the final generated trajectories dataset $\tilde{\mathcal{D}}$ for offline RL training.
In summary, our BiTrajDiff operates in two key stages. Initially, it trains a bidirectional diffusion model to capture the distributions of both forward-future and backward-history trajectories. Subsequently, BiTrajDiff iteratively generates bidirectional trajectory conditioned on intermediate anchor states sampled from $\mathcal{D}$. These generated sequences undergo further completion and filtering to create a high-fidelity augmented dataset $\tilde{\mathcal{D}}$, which is then mixed with $\mathcal{D}$ to train downstream offline RL algorithms. We adopt CleanDiffuser (dong2024cleandiffuser) as the backbone for BiTrajDiff, owing to its generality and robustness. The pseudocode of BiTrajDiff is presented in Appendix A.
@@ 5 Experiments
To demonstrate the effectiveness of the proposed BiTrajDiff method, we conduct experiments on the D4RL benchmark (fu2020d4rl). Our evaluation seeks to answer the following questions: (1) Does BiTrajDiff outperform existing DA methods across various offline RL algorithms? (Section 5.2 and Appendix C) (2) How do different components of BiTrajDiff affect the offline RL performance? (Section 5.3 and Appendix D-F) (3) Can BiTrajDiff find the potentially valuable state leading to high return more effectively than other DA baselines? (Section 5.4 and Appendix G) (4) Can BiTrajDiff generate more diverse trajectories compared to single-direction DA approaches? (Section 5.5)
@@ 5.1 Experiment Settings
We evaluate our approach on the D4RL benchmark (fu2020d4rl), focusing on locomotion, navigation, and manipulation tasks. We compare BiTrajDiff with three state-of-the-art DA methods: Synther (lu2023synthetic),DiffStitch (li2024diffstitch) and RTDiff (yang2025rtdiff). To evaluate the effectiveness of them, we conduct extensive experiments on four representative offline RL algorithms: IQL (kostrikov2021offline), TD3BC (fujimoto2021minimalist), CQL (kumar2020conservative), and DT (chen2021decision). These methods have been widely adopted as standard baselines in offline RL due to their stable performance. The hyperparameter setting can be viewed in Appendix B.2.
Table 1: Performance of our BiTrajDiff and baselines on the locomotion tasks. $\pm$ corresponds to the standard deviation of the performance on 5 random seeds. The best and the second-best results of each setting are marked as bold and underline, respectively. Detailed reports about CQL and DT are reported in Appendix C.
Source Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)